OpenStudy (zmudz):

Find a closed form for \(S_n = 1 \cdot 1! + 2 \cdot 2! + \ldots + n \cdot n!.\) for integer \(n \geq 1.\) Your response should have a factorial.

ganeshie8 (ganeshie8):

\[\begin{align}S_n &= \color{red}{1} \cdot 1! + \color{red}{2} \cdot 2! + \ldots + \color{red}{n} \cdot n!\\~\\ &= \color{red}{(2-1)} \cdot 1! + \color{red}{(3-1)} \cdot 2! + \ldots + \color{red}{(n+1-1)} \cdot n!\\~\\ &=2!-1! + 3!-2! +\cdots + (n+1)!-n!\\~\\ &=-1! + (n+1)! \end{align}\]

OpenStudy (anonymous):

You might recall asking this about a month ago... http://openstudy.com/users/zmudz#/updates/55ad142ee4b0d48ca8ed35d0